# Searching for quantum-gravity footprint around stellar-mass black holes

### v2

@@ 4.6 GRO J0422+32
The latest object is a candidate for the smallest black hole known, with $M=2.7_{-0.5}^{+0.7}M_{\odot}$, $d=2.49\pm 0.30$ kpc, $i=55.6\pm 4.1$ deg (Gelino & Harrison, 2003; Casares et al., 2022). No information about the spin was found, and therefore it is not possible to estimate the expected $r_{\rm isco}$. We found only one paper useful to estimate the inner radius: $r_{\rm in}=(5.1\pm 2.3)r_{\rm g}$ (Shrader et al., 1997). We also found one Swift observation (id 00032976001) performed on October 3, 2013. We analyzed the data with the same procedure adopted for the above case, but we did not find any source, with an upper limit of the flux of $\sim 1.5\times 10^{-11}$ erg cm-2 s-1 ($3\sigma$ detection, $\sim 1.2$ ks exposure in photon counting mode, $N_{\rm H}=1.51\times 10^{21}$ cm-2 HI4PI Collaboration (2016), power-law model with $\Gamma=2$). Also GRO J0422+32 was likely in a quiescent state.
@@ 5 Conclusions
We searched in all the available literature for observations of X-ray binaries in high/soft state, where it is expected that the inner disk is closest to the innermost stable circular orbit ($r_{\rm in}\sim r_{\rm isco}$) according to the Kerr metric, to test if there are deviations from general relativity. We selected the measurements made by using the diskbb model, we applied the best corrections for the Comptonization (Shimura & Takahara, 1995) and the inner boundary (Kubota et al., 1998), we made all the values homogeneous according to the most recent and reliable values of the reference quantities ($r_{\rm g}$, $a$, $d$, $i$). All measured radii are consistent with the expectations of general relativity. Some anomalous cases can be easily reconciled by taking into account Comptonization and how it affects the estimate of $T_{\rm in}$, uncertainties in the reference quantities, and instrumental biases.
@@
$\tilde{\xi}=\frac{\xi}{r_{\rm g}^{2}}\gtrsim 0.028$ (9)
As expected, given the currently available instruments, a significant weakening of $G$ is needed to obtain a measurable reduction of $r_{\rm isco}$, according to the Sánchez (2024) theory. If NewAthena [^7] https://www.cosmos.esa.int/web/athena/home is realized, it will be possible to improve the statistics significantly.
We note that a change in $G$ would affect also the measurement of the mass of the black hole. In the case of XRBs, the mass is measured via the Kepler’s third law, which includes $G$ (Casares & Jonker, 2014). Therefore, we should revise the mass of Cyg X-1 by using Eq. (1) in the mass function equation (cf Eq. 1 in Casares & Jonker 2014). However, the orbit of the binary system (the semimajor axis is $\sim 0.244\,\mathrm{AU}\sim 3.7\times 10^{7}$ km, Miller-Jones et al. 2021) is much greater than the size of $r_{\rm isco}=1.61r_{\rm g}\sim 50$ km, and, given a dependence on $r^{2}$ of $G(r)$ (cf Eq. 1), the effect is negligible and smaller than the present measurement error.
We cannot set any constraint on negative values of $\xi$ with this method. In this case, $r_{\rm isco}$ from ASG should be greater than the value from GR. However, a greater $r_{\rm in}$ is expected from state transitions in stellar-mass black holes. Particularly, in the hard state, the inner disk is truncated and $r_{\rm in}$ is of the order of tens $r_{\rm g}$ (see, however, Datta et al. 2024).
Other, more complex expressions of $G$ are currently under study (Bonanno et al., 2024), and will be the topic of another essay.
@@
Acknowledgments
LF thanks Roberto Della Ceca, director of the Brera Astronomical Observatory – INAF, for partially funding this research, and Daniele Malafarina for useful questions after the presentation of this work at the IV International FLAG workshop: the Quantum and the Gravity, (Catania, Italy, 9-11 September 2024). We thank the anonymous reviewers, whose comments helped to improve the essay.

### v3

@@ 4.6 GRO J0422+32
The latest object is a candidate for the smallest black hole known, with $M=2.7_{-0.5}^{+0.7}M_{\odot}$, $d=2.49\pm 0.30$ kpc, $i=55.6\pm 4.1$ deg [35, 16]. No information about the spin was found, and therefore it is not possible to estimate the expected $r_{\rm isco}$. We found only one paper useful to estimate the inner radius: $r_{\rm in}=(5.1\pm 2.3)r_{\rm g}$ [100]. We also found one Swift observation (id 00032976001) performed on October 3, 2013. We analyzed the data with the same procedure adopted for the above case, but we did not find any source, with an upper limit of the flux of $\sim 1.5\times 10^{-11}$ erg cm-2 s-1 ($3\sigma$ detection, $\sim 1.2$ ks exposure in photon counting mode, $N_{\rm H}=1.51\times 10^{21}$ cm-2 [38], power-law model with $\Gamma=2$). Also GRO J0422+32 was likely in a quiescent state.
@@ 5 Conclusions
We searched in all the available literature for observations of X-ray binaries in high/soft state, where it is expected that the inner disk is closest to the innermost stable circular orbit ($r_{\rm in}\sim r_{\rm isco}$) according to the Kerr metric, to test if there are deviations from general relativity. We selected the measurements made by using the diskbb model, we applied the best corrections for the Comptonization [98] and the inner boundary [42], we made all the values homogeneous according to the most recent and reliable values of the reference quantities ($r_{\rm g}$, $a$, $d$, ii). All measured radii are consistent with the expectations of general relativity. Some anomalous cases can be easily reconciled by taking into account Comptonization and how it affects the estimate of $T_{\rm in}$, uncertainties in the reference quantities, and instrumental biases.
@@
$\tilde{\xi}=\frac{\xi}{r_{\rm g}^{2}}\gtrsim 0.24$ (9)
corresponding to $\xi>2.3\times 10^{12}$ cm2. As expected, given the currently available instruments, a significant weakening of $G$ is needed to obtain a measurable reduction of $r_{\rm isco}$, according to the Sánchez’s [95] theory. Nonetheless, this is the tightest constraint obtained to date. If NewAthena [^8] https://www.cosmos.esa.int/web/athena/home will be realized, it will be possible to improve the statistics significantly.
We note that a change in $G$ would affect also the measurement of the mass of the black hole. In the case of XRBs, the mass is measured via the Kepler’s third law, which includes $G$ [15]. Therefore, we should revise the mass of Cyg X-1 by using Eq. (1) in the mass function equation (cf Eq. 1 in [15]). However, the orbit of the binary system (the semimajor axis is $\sim 0.244\,\mathrm{AU}\sim 3.7\times 10^{7}$ km, [63]) is much greater than the size of $r_{\rm isco}=1.61r_{\rm g}\sim 50$ km, and, given a dependence on $r^{2}$ of $G(r)$ (cf Eq. 1), the effect is negligible and smaller than the present measurement error.
We cannot set any constraint on negative values of $\xi$ with this method. In this case, $r_{\rm isco}$ from ASG should be greater than the value from GR. However, a greater $r_{\rm in}$ is expected from state transitions in stellar-mass black holes. Particularly, in the hard state, the inner disk is truncated and $r_{\rm in}$ is of the order of tens $r_{\rm g}$ (see, however, [20]).
Other, more complex expressions of $G$ are currently under study [13], and will be the topic of another essay.
@@
[…]
Acknowledgements LF thanks Roberto Della Ceca, director of the Brera Astronomical Observatory – INAF, for partially funding this research, and Daniele Malafarina for useful questions after the presentation of this work at the IV International FLAG workshop: the Quantum and the Gravity, (Catania, Italy, 9-11 September 2024).